### v1

@@ 3.3 Vorticity scaling
We observe that the intensity of vorticity fluctuations, $\langle\omega_{\alpha}^{\prime}\omega_{\alpha}^{\prime}\rangle^{1/2}$ (no summation over Greek indices), in OCF collapses well with corresponding closed channel flow profiles over most of the channel height, except for the region very close to the free surface (figure omitted). The profile of the wall-normal vorticity component in OCF already starts to deviate from the CCF profile approximately at the bottom of the normal velocity damping layer $\delta_{NVD}$.
@@
In Antonia et al. [2] and Panton [34] it was shown that the vorticity fluctuation intensity in the outer layer of wall-bounded turbulent flows scales with the Kolmogorov time scale, $\tau_{\eta}=(\nu/\varepsilon)^{1/2}=(\nu h/u_{\tau}^{3})^{1/2}$ (assuming that the dissipation rate $\varepsilon$ scales with $u_{\tau}$ and $h$), leading to $\omega_{i,rms}{\sim}\sqrt{(\nu h)/u^{3}_{\tau}}$. CM03 found the same scaling to be true for the vorticity fluctuation intensity in open channel flow up to the free surface and, thus, suggested to multiply the vorticity fluctuation intensity with $\mathrm{Re}_{\tau}^{1/2}$ after being normalized in wall units in order to collapse data at different Reynolds numbers. Applying the above normalization in the upper part of the channel domain, we indeed achieve a satisfactory collapse of the different Reynolds number profiles away from the wall, as shown in figure 9. It can be seen that the channel height $h$ is the characteristic length scale away both from the wall and the free surface (cf. figure 9a).
Figure 9: Vorticity fluctuation intensity near the free surface, normalized by $u_{\tau}^{2}/\nu\mathrm{Re}_{\tau}^{-1/2}$. , $\omega_{x,rms}$; , $\omega_{y,rms}$; , $\omega_{z,rms}$; solid lines, O200; dashed lines, O400, dashed dotted lines, O600; dotted lines, O900. The distance from the free surface is normalized with (a) the channel height $h$; (b) the near-surface viscous length scale $\ell_{V}$; (c) the near-surface Kolmogorov length scale $\ell_{K}$.
@@
Figure 10: Conditional average of vorticity fluctuation intensity near the free surface, normalized by $u_{\tau}^{2}/\nu\mathrm{Re}_{\tau}^{-1/2}$. Legend see figure 9. (a,b,c) $u^{\prime}>0$; (d,e,f) $u^{\prime}<0$. The distance from the free surface is normalized with (a,d) the channel height $h$; (b,e) the near-surface viscous length scale $\ell_{V}$; (c,f) the near-surface Kolmogorov length scale $\ell_{K}$.
@@ 3.4 Super-streamwise vortices (SSV)
In order to extract SSVs from the DNS data, we compute the streamfunction of the streamwise-averaged crossflow velocity components, defined as follows:
$\psi_{\langle v\rangle_{x}\langle w\rangle_{x}}(y,z,t)=\int_{\hat{y}}^{y}\langle w\rangle_{x}(\hat{y},z,t)\,{\rm d}\hat{y}-\int_{\hat{z}}^{z}\langle v\rangle_{x}(y,\hat{z},t)\,{\rm d}\hat{z}\,.$ (6)
@@ 3.5 Updated surface-influenced layer
The comparison of the streamwise turbulence intensity between OCF and CCF in section 3.2 and SSV in section 3.4 revealed that the influence of the free-slip boundary condition reaches much further towards the solid wall than previously assumed. In order to account for this new evidence, we separate the “surface-influenced layer” from a new “normal-velocity-damping layer”, thereby expanding the previous three-layer structure to four layers, as sketched in figure 13. The thickness of the former layer is found to extend essentially across the entire channel, hence we propose $\delta_{SIL}\approx h$. The thickness of the latter layer, where the surface-normal turbulence intensity profile in OCF differs from CCF, is quantified by $\delta_{NVD}\approx 0.3h$, similar to previous studies [e.g. 29, 12]. On the other hand, the near-surface viscous sublayer, where the surface-normal velocity grows linearly when moving away from the free surface, measures $\delta_{V}\approx\ell_{V}$, and the Kolmogorov sublayer, where the surface-normal gradient of the mean velocity as well as the intensities of the fluctuations of the surface-parallel vorticity components are damped to zero, is estimated as $\delta_{K}\approx 20\ell_{K}$.
Figure 13: Updated surface-influenced layer structure in OCF for high Reynolds numbers. Surface-influenced layer $\delta_{SIL}\approx h$, normal-velocity-damping layer $\delta_{NVD}\approx 0.3h$, near-surface viscous sublayer $\delta_{V}\approx\ell_{V}$, Kolmogorov sublayer $\delta_{K}\approx 20\ell_{K}$.
@@
Table 2: Surface-influenced layer widths ($\delta_{SIL}$, $\delta_{NVD}$, $\delta_{V}$, $\delta_{K}$).
case $\mathrm{Re}_{\tau}$ $\delta_{SIL}/h$ $\delta_{NVD}/h$ $\delta_{V}/h$ $\delta_{K}/h$ $\delta_{V}^{+}$ $\delta_{K}^{+}$
@@
Note that $\delta_{K}>\delta_{V}$ in the present range of Reynolds numbers, while Reynolds numbers $\mathrm{Re}_{b}\gtrsim 160000$ are necessary for $\delta_{K}<\delta_{V}$. Hence the relative widths of $\delta_{V}$ and $\delta_{K}$ sketched in figure 13 apply to high-Reynolds number OCF.
The Kolmogorov layer width, which varies only little when normalized in wall units for the range of Reynolds numbers considered ($14.2\leq\delta_{K}^{+}\leq 17.6$), is of the same order of magnitude as the wall-normal extent of the viscous sublayer adjacent to the no-slip wall. This highlights a numerical resolution requirement for the free-slip boundary similar to the one near the no-slip wall. DNSs of OCF often use one-sided grid refinement only, which leads to acceptable one-point turbulence statistics near the solid wall. However, the non-linear interactions between the small scales, which are located in the near-surface region, and larger scales of motions might not be captured correctly without additional grid refinement towards the free-slip boundary.
@@ 4 Conclusion
In turbulent open channel flow the free surface is typically modelled as a free-slip boundary when it can be expected to undergo only small deformations. As a consequence the statistics generally differ from the counterpart in closed channel flow. For a long time the differences in these two configurations were believed to be restricted to the vicinity of the free surface. The most prominent effect of the free-slip boundary condition is on the structure of the Reynolds stress tensor, which has been widely discussed in the literature [39, 17, 35, 32, 29, 16, 8, 30, 31]. In this work, the scaling of turbulence statistics in the vicinity of a free-slip boundary in open channel flow has been analyzed with the aid of data from DNS in computational domains large enough to capture the largest turbulence scales while providing sufficient grid refinement near the free-slip boundary. For $\mathrm{Re}_{\tau}$ up to 900, the triple surface layer proposed by CM03 has been observed for the spanwise and wall-normal Reynolds stress components as well as for the vorticity fluctuation intensities. In accordance with CM03 we found the channel height $h$, the viscous length $\ell_{V}$, and the Kolmogorov length $\ell_{K}$ to be the relevant length scales for the near-surface layers. Unlike CM03, however, the sublayer where the wall-normal velocity decreases linearly towards the surface is found to scale with the near-surface viscous length scale $\ell_{V}$ (instead of $\ell_{K}$), and the thin near-surface vorticity layer is found to scale with the Kolmogorov length scale $\ell_{K}$ (instead of $\ell_{V}$). Regarding their intensity, spanwise and wall-normal velocity fluctuations do scale in wall units, while the streamwise component tends to increase under that scaling when increasing the Reynolds number. This failure of wall-scaling is well known for closed channel flow and it is related to VLSMs that contribute to the streamwise Reynolds stress component for higher Reynolds numbers. For open channel flows this effect appears already at lower Reynolds numbers and leads—unlike for closed channel flows—to a collapse of the streamwise turbulent intensity profiles normalized with the bulk velocity in the region between $y/h\approx 0.3$ and $y/h=1$, which is already visible for $\mathrm{Re}_{\tau}{\geq}400$. Since the footprint of the enhanced VLSMs can be found in energy spectra very close to the solid wall, the so-called surface-influenced layer $\delta_{SIL}$, described by CM03 and others, has to be extended essentially all the way to the solid boundary for OCF. As a consequence, turbulence in the vicinity of a free-slip boundary in OCF has effectively to be characterized by four layers: The Kolmogorov sublayer $\delta_{K}\approx 20\mathrm{Re}_{b}^{-3/4}h$—where surface-parallel vorticities are damped to zero—, the viscous sublayer $\delta_{V}\approx\mathrm{Re}_{b}^{-1/2}h$—where the surface-normal velocity component decreases linearly to zero—, the normal velocity damping layer $\delta_{NVD}\approx 0.3h$—where the surface-normal velocity profile deviates from the CCF profile— and the surface-influenced layer spanning the entire channel height $\delta_{SIL}=h$—where the streamwise turbulence intensity profile deviates from the CCF profile.
@@
Acknowledgements Many fruitful discussions with Genta Kawahara throughout this work are gratefully acknowledged. The simulations were carried out at the HPC clusters UC2 and CARA. Thus, the authors gratefully acknowledge the scientific support and HPC resources provided by the SCC Karlstuhe, the German Aerospace Center (DLR), and the state of Baden-Wuerttemberg through bwHPC. The HPC system CARA is partially funded by ”Saxon State Ministry for Economic Affairs, Labour and Transport” and ”Federal Ministry for Economic Affairs and Climate Action”. The authors acknowledge funding by DFG through grant UH 242/3-1.

### v2

@@ 3.3 Vorticity scaling
We observe that the intensity of vorticity fluctuations, $\langle\omega_{\alpha}^{\prime}\omega_{\alpha}^{\prime}\rangle^{1/2}$ (no summation over Greek indices), in OCF collapses well with corresponding closed channel flow profiles over most of the channel height, except for the region very close to the free surface (figure omitted). The profile of the wall-normal vorticity component in OCF already starts to deviate from the CCF profile approximately at the bottom of the normal velocity damping layer $\delta_{\mathrm{NVD}}$.
@@
In Antonia et al. [2] and Panton [34] it was shown that the vorticity fluctuation intensity in the outer layer of wall-bounded turbulent flows scales with the Kolmogorov time scale, $\tau_{\eta}=(\nu/\varepsilon)^{1/2}=(\nu h/u_{\tau}^{3})^{1/2}$ (assuming that the dissipation rate $\varepsilon$ scales with $u_{\tau}$ and $h$), leading to $\omega_{i,rms}{\sim}\sqrt{u^{3}_{\tau}/(\nu h)}$. CM03 found the same scaling to be true for the vorticity fluctuation intensity in open channel flow up to the free surface and, thus, suggested to multiply the vorticity fluctuation intensity with $\mathrm{Re}_{\tau}^{1/2}$ after being normalized in wall units in order to collapse data at different Reynolds numbers. Applying the above normalization in the upper part of the channel domain, we indeed achieve a satisfactory collapse of the different Reynolds number profiles away from the wall, as shown in figure 9. It can be seen that the channel height $h$ is the characteristic length scale away both from the wall and the free surface (cf. figure 9a).
Figure 9: Vorticity fluctuation intensity near the free surface, normalized by $u_{\tau}^{2}/\nu\mathrm{Re}_{\tau}^{-1/2}$. , $\omega_{x,rms}$; , $\omega_{y,rms}$; , $\omega_{z,rms}$; solid lines, O200; dashed lines, O400, dashed dotted lines, O600; dotted lines, O900. The distance from the free surface is normalized with (a) the channel height $h$; (b) the near-surface viscous length scale $\ell_{\mathrm{V}}$; (c) the near-surface Kolmogorov length scale $\ell_{\mathrm{K}}$.
@@
Figure 10: Conditional average of vorticity fluctuation intensity near the free surface, normalized by $u_{\tau}^{2}/\nu\mathrm{Re}_{\tau}^{-1/2}$. Legend see figure 9. (a,b,c) $u^{\prime}>0$; (d,e,f) $u^{\prime}<0$. The distance from the free surface is normalized with (a,d) the channel height $h$; (b,e) the near-surface viscous length scale $\ell_{\mathrm{V}}$; (c,f) the near-surface Kolmogorov length scale $\ell_{\mathrm{K}}$.
@@ 3.4 Super-streamwise vortices (SSV)
In order to extract SSVs from the DNS data, we compute the streamfunction of the streamwise-averaged crossflow velocity components, $\psi_{\langle v\rangle_{x}\langle w\rangle_{x}}(y,z,t)$, that fulfills the requirement
$\begin{pmatrix}\partial\psi_{\langle v\rangle_{x}\langle w\rangle_{x}}/\partial y\\ \partial\psi_{\langle v\rangle_{x}\langle w\rangle_{x}}/\partial z\end{pmatrix}=\begin{pmatrix}-\langle w\rangle_{x}\\ \langle v\rangle_{x}\end{pmatrix}\,.$ (6)
@@ 3.5 Updated surface-influenced layer
The comparison of the streamwise turbulence intensity between OCF and CCF in section 3.2 and SSV in section 3.4 revealed that the influence of the free-slip boundary condition reaches much further towards the solid wall than previously assumed. In order to account for this new evidence, we separate the “surface-influenced layer” from a new “normal-velocity-damping layer”, thereby expanding the previous three-layer structure to four layers, as sketched in figure 13. The thickness of the former layer is found to extend essentially across the entire channel, hence we propose $\delta_{\mathrm{SIL}}\approx h$. The thickness of the latter layer, where the surface-normal turbulence intensity profile in OCF differs from CCF, is quantified by $\delta_{\mathrm{NVD}}\approx 0.3h$, similar to previous studies [e.g. 29, 12]. On the other hand, the near-surface viscous sublayer, where the surface-normal velocity grows linearly when moving away from the free surface, measures $\delta_{\mathrm{V}}\approx\ell_{\mathrm{V}}$, and the Kolmogorov sublayer, where the surface-normal gradient of the mean velocity as well as the intensities of the fluctuations of the surface-parallel vorticity components are damped to zero, is estimated as $\delta_{\mathrm{K}}\approx 20\ell_{\mathrm{K}}$.
Figure 13: Updated surface-influenced layer structure in OCF for high Reynolds numbers. Surface-influenced layer $\delta_{\mathrm{SIL}}\approx h$, normal-velocity-damping layer $\delta_{\mathrm{NVD}}\approx 0.3h$, near-surface viscous sublayer $\delta_{\mathrm{V}}\approx\ell_{\mathrm{V}}$, Kolmogorov sublayer $\delta_{\mathrm{K}}\approx 20\ell_{\mathrm{K}}$.
@@
Table 2: Surface-influenced layer widths ($\delta_{\mathrm{SIL}}$, $\delta_{\mathrm{NVD}}$, $\delta_{\mathrm{V}}$, $\delta_{\mathrm{K}}$).
case $\mathrm{Re}_{\tau}$ $\delta_{\mathrm{SIL}}/h$ $\delta_{\mathrm{NVD}}/h$ $\delta_{\mathrm{V}}/h$ $\delta_{\mathrm{K}}/h$ $\delta_{\mathrm{V}}^{+}$ $\delta_{\mathrm{K}}^{+}$
@@
Note that $\delta_{\mathrm{K}}>\delta_{\mathrm{V}}$ in the present range of Reynolds numbers, while Reynolds numbers $\mathrm{Re}_{\mathrm{b}}\gtrsim 160000$ are necessary for $\delta_{\mathrm{K}}<\delta_{\mathrm{V}}$. Hence the relative widths of $\delta_{\mathrm{V}}$ and $\delta_{\mathrm{K}}$ sketched in figure 13 apply to high-Reynolds number OCF.
The Kolmogorov layer width, which varies only little when normalized in wall units for the range of Reynolds numbers considered ($14.2\leq\delta_{\mathrm{K}}^{+}\leq 17.6$), is of the same order of magnitude as the wall-normal extent of the viscous sublayer adjacent to the no-slip wall. This highlights a numerical resolution requirement for the free-slip boundary similar to the one near the no-slip wall. DNSs of OCF often use one-sided grid refinement only, which leads to acceptable one-point turbulence statistics near the solid wall. However, the non-linear interactions between the small scales, which are located in the near-surface region, and larger scales of motions might not be captured correctly without additional grid refinement towards the free-slip boundary.
@@ 4 Conclusion
In turbulent open channel flow the free surface is typically modelled as a free-slip boundary when it can be expected to undergo only small deformations. As a consequence the statistics generally differ from the counterpart in closed channel flow. For a long time the differences in these two configurations were believed to be restricted to the vicinity of the free surface. The most prominent effect of the free-slip boundary condition is on the structure of the Reynolds stress tensor, which has been widely discussed in the literature [39, 17, 35, 32, 29, 16, 8, 30, 31]. In this work, the scaling of turbulence statistics in the vicinity of a free-slip boundary in open channel flow has been analyzed with the aid of data from DNS in computational domains large enough to capture the largest turbulence scales while providing sufficient grid refinement near the free-slip boundary. For $\mathrm{Re}_{\tau}$ up to 900, the triple surface layer proposed by CM03 has been observed for the spanwise and wall-normal Reynolds stress components as well as for the vorticity fluctuation intensities. In accordance with CM03 we found the channel height $h$, the viscous length $\ell_{\mathrm{V}}$, and the Kolmogorov length $\ell_{\mathrm{K}}$ to be the relevant length scales for the near-surface layers. Unlike CM03, however, the sublayer where the wall-normal velocity decreases linearly towards the surface is found to scale with the near-surface viscous length scale $\ell_{\mathrm{V}}$ (instead of $\ell_{\mathrm{K}}$), and the thin near-surface vorticity layer is found to scale with the Kolmogorov length scale $\ell_{\mathrm{K}}$ (instead of $\ell_{\mathrm{V}}$). Regarding their intensity, spanwise and wall-normal velocity fluctuations do scale in wall units, while the streamwise component tends to increase under that scaling when increasing the Reynolds number. This failure of wall-scaling is well known for closed channel flow and it is related to VLSMs that contribute to the streamwise Reynolds stress component for higher Reynolds numbers. For open channel flows this effect appears already at lower Reynolds numbers and leads—unlike for closed channel flows—to a collapse of the streamwise turbulent intensity profiles normalized with the bulk velocity in the region between $y/h\approx 0.3$ and $y/h=1$, which is already visible for $\mathrm{Re}_{\tau}{\geq}400$. Since the footprint of the enhanced VLSMs can be found in energy spectra very close to the solid wall, the so-called surface-influenced layer $\delta_{\mathrm{SIL}}$, described by CM03 and others, has to be extended essentially all the way to the solid boundary for OCF. As a consequence, turbulence in the vicinity of a free-slip boundary in OCF has effectively to be characterized by four layers: The Kolmogorov sublayer $\delta_{\mathrm{K}}\approx 20\mathrm{Re}_{\mathrm{b}}^{-3/4}h$—where surface-parallel vorticities are damped to zero—, the viscous sublayer $\delta_{\mathrm{V}}\approx\mathrm{Re}_{\mathrm{b}}^{-1/2}h$—where the surface-normal velocity component decreases linearly to zero—, the normal velocity damping layer $\delta_{\mathrm{NVD}}\approx 0.3h$—where the surface-normal velocity profile deviates from the CCF profile— and the surface-influenced layer spanning the entire channel height $\delta_{\mathrm{SIL}}=h$—where the streamwise turbulence intensity profile deviates from the CCF profile.
@@
Acknowledgements Many fruitful discussions with Genta Kawahara throughout this work are gratefully acknowledged. The simulations were carried out at the HPC clusters UC2 and CARA. Thus, the authors gratefully acknowledge the scientific support and HPC resources provided by the SCC Karlstuhe, the German Aerospace Center (DLR), and the state of Baden-Wuerttemberg through bwHPC. The HPC system CARA is partially funded by ”Saxon State Ministry for Economic Affairs, Labour and Transport” and ”Federal Ministry for Economic Affairs and Climate Action”.
Declarations
  • Funding: The authors acknowledge funding by DFG through grant UH 242/3-1.